A company prices is hurricane insurance using the following assumptions
i. In any calendar year, there can be at most one hurricane
ii. In any calendar year,the probability of a hurricane is 0.05
iii. The number of hurricanes in any calendar year is independent of the number of hurricanes in any other calendar year.
Using the company’s assumptions, calculate the probability that there are fewer than 3 hurricanes in a 20 year period.
1
Expert's answer
2020-05-12T19:11:17-0400

By the binomial distribution, this probability is

P(X<3)=P(x=0)+P(x=1)+P(x=2)

P(x=0)=C20,0(0.05)0(1-0.05)20-0=C20,0(0.05)0(0.95)20=0.358

P(x=1)=C20,1(0.05)1(0.95)19=0.377

P(x=2)=C20,2(0.05)2(0.95)18=0.189

P=0.358+0.377+0.189=0.924